OpenStudy (anonymous):

what is the coefficient of x in the division (18x^3+12x^2-3x) divided by 6x^2

OpenStudy (anonymous):

2. divide (2x^3-x^2-24x+12) divided by (2x-1 3.simplify 2a+2/a^2-1 4. what is the value of b? 3b^2+13b+4/b+4 5.multiply x^2 +8x+15/x-4 times x^2-16/2x+6 9. solve 10/3x+4/3=7+x/2x 10. solve x+12/x+4=x/x+8 (im not good at word probs) 11. John walked 10 miless on saturday. he walked twice as fast on the second 5miles f his walk than walked on the first 5miles of his walk. Write an expression that representss the time he spent walking...Let x=John's speed onthe first half of his walk. THANK YOU FOR THE HELP I HOPE TO GET>>>>ALGEBRA IS NOT AT ALL 1 OF MY BEST SUBJECTS!!! O_o

OpenStudy (anonymous):

never mind n 11. i got it